# LayerCollapse: Adaptive compression of neural networks

### v1

@@ Appendix A Proof of Equation (4)
Appendix B Experiment Details
@@
B.1 Stand-Alone Compression
@@
Training Setup: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques. The loss function combined cross-entropy and KL Divergence terms:
@@
B.2 LayerCollapse vs. Knowledge Distillation
For both LayerCollapse and Knowledge Distillation, the target architectures were identical. The LayerCollapse setup followed the procedure outlined in Section B.1. The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau. Experiments concluded when no further progress was achievable.
B.3 Ablation Study of Regularization
@@
B.4 LayerCollapse in NLP
We utilized the base Bert model from Hugging Face for NLP experiments. Finetuning employed the Adam optimizer with a learning rate of $0.00002$.

### v3

@@ Appendix A Proof of Equation (4)
Appendix B AdamLC
Inspired by AdamW Loshchilov & Hutter (2017) we decouple the adam optimizer with our regularization so the adaptive learning rate does not make the regularization un-predictable the algorithm is given in 2. In this algorithm $lc$ is the regularization constant.
Algorithm 2 AdamLC Optimization Algorithm
0:  $\alpha,\beta_{1},\beta_{2},\epsilon,\theta_{0},lc$
0:  Optimized parameters $\theta$
1:  Initialize time step $t\leftarrow 0$
2:  Initialize first moment vector $m_{0}\leftarrow\mathbf{0}$
3:  Initialize second moment vector $v_{0}\leftarrow\mathbf{0}$
4:  Initialize parameter vector $\theta\leftarrow\theta_{0}$
5:  while not converged do
6:     $t\leftarrow t+1$
7:     $g_{t}\leftarrow\text{gradient at }\theta$
8:     $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}+(1-\beta_{1})\cdot g_{t}$
9:     $v_{t}\leftarrow\beta_{2}\cdot v_{t-1}+(1-\beta_{2})\cdot g_{t}^{2}$
10:     $\hat{m}_{t}\leftarrow\frac{m_{t}}{1-\beta_{1}^{t}}$
11:     $\hat{v}_{t}\leftarrow\frac{v_{t}}{1-\beta_{2}^{t}}$
12:     $\theta\leftarrow\theta-\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
13:     if $\theta$ has a negative slope then
14:        $\theta\leftarrow\theta-2\cdot(\theta-1)\cdot lc\cdot\alpha$
15:     end if
16:  end while
17:  return $\theta$
Appendix C Experiment Details
@@
C.1 Regularization Performance of LayerCollapse
Training was conducted using a modified version of the Adam optimizer, as detailed in Appendix B. A consistent regularization parameter, Regularization 5, was applied across all experiments, and the same value was used for comparisons with the LayerDrop method. The learning rate scheduler LiveTune Shabgahi et al. (2023) was employed, with an initial learning rate set to $0.001$.
In all experiments, a batch size of 128 was maintained, and the models were trained for a total of 3,000 steps. For performance evaluation, we applied a stride of 32 on the complete test dataset of Wikitext-103.
C.2 LayerCollapse for Compression
C.2.1 Performance on GLUE
Because of the diversity of tasks each have their own requirements for training (i.e. optimizer, learning rate, and epochs) All tasks except for QQP the SGD optimizer is used, and the modified adamLC B is used for QQP. In all experiments initial learning rate is 0.0001 which reduces to 0.00001 linearly. Collapsing is done one layer at a time. Starting from the final un-collapsed layer. In each step we reduce the regularization loss of the active layer during a retraining phase, when the loss is bellow $1e^{-4}$ we move to the next layer. All tasks for Bert-base use batch size 128. Bert-large is trained on batch size 20.
C.2.2 Performance of ImageNet-1k
@@
Self Distillation: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques. The loss function combined cross-entropy and KL Divergence terms:
@@
C.3 LayerCollapse vs. Knowledge Distillation
For both LayerCollapse and Knowledge Distillation, the target architectures were identical. The LayerCollapse setup followed the procedure outlined in Section C.2.2. The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau, and concluded when no further progress was possible.
C.4 Study of Regularization on VGG